Simplifying Square Roots
How to simply these square roots;
1. 25
2. 3 * 7
3. 48/27
4. 300

Answer by checkley71(8403)

sqrt25=sqrt5*5=5 answer.
sqrt3*7=sqrt21 answer.
sqrt48/27=sqrt(16*3)/(9*3)=4/3 answer.
sqrt300=sqrt(100*3)=10sqrt3 answer.